# Hardware-Compatible Single-Shot Feasible-Space Heuristics for Solving the Quadratic Assignment Problem

### v1

@@ 5.4 Computation Breakdown of the Parallel Search and Its Applications to IMC Hardware
In this section, we present a computational breakdown of our Python-based CPU implementation of Algorithm 4.5 to identify the computationally intensive routines by varying the problem size from $n=12$ to $n=100$. Specifically, we compare the computation load distribution of the subroutines in Algorithm 4.5 relative to the total runtime. The instance names are given in each pie chart in Figure 5.6, followed by the average runtime for $10$ repetitions in terms of the wall-clock time. The routines for computing $g$ and $\tilde{\Delta}$ (line 14 in Algorithm 4.5) are labelled “$g$ computation” and “$\tilde{\Delta}$ computation”, respectively. The update of the full neighbourhood (line 8 in Algorithm 4.5) is labelled “Neighbourhood $\mathcal{N}$ update”, and all other routines are collectively labelled “Other”. The comparison is Figure 5.6 was generated using a MacBook Pro equipped with an Apple M3 chip and 16 GB of memory. The Python profiler named “cProfile” was used to provide the performance breakdown.
@@
In Figure 5.6, we observe a clear trend in which the fraction of the computational time required for calculating $g$ and $\tilde{\Delta}$ increases as the problem size grows. This increase in time is due primarily to the fact that these operations involve matrix–vector multiplications. As the problem size increases, the matrix–vector multiplications become more demanding due to the quadratic growth of the binary formulation. Given that these steps are computational bottlenecks, IMC hardware presents a promising opportunity for considerable improvements in terms of computation time. Unlike conventional processors, which suffer from frequent memory access overhead, IMC architectures perform computations directly within memory. As shown in Figure 4.3, our approach is well-suited for this paradigm. Furthermore, the most demanding computation shown in Figure 5.6 is the $\tilde{\Delta}$ computation, which consists of a simple matrix–vector multiplication. Additionally, the energy efficiency of IMC hardware is a notable by-product, as it reduces the energy overhead of conventional digital computers.

### v2

@@ 5.3 Computation Breakdown of the Parallel Search and Its Applications to IMC Hardware
In this section, we present a computational breakdown of our Python-based CPU implementation of Algorithm 4.5 to identify the computationally intensive routines by varying the problem size from $n=12$ to $n=100$. Specifically, we compare the computation load distribution of the subroutines in Algorithm 4.5 relative to the total runtime. The instance names are given in each pie chart in Figure 5.7, followed by the average runtime for $10$ repetitions in terms of the wall-clock time. The routines for computing $g$ and $\tilde{\Delta}$ (line 14 in Algorithm 4.5) are labelled “Gradient computation”. The update of the full neighbourhood (line 8 in Algorithm 4.5) is labelled “Neighbourhood $\mathcal{N}$ update”. The routine for implementing the Top10 heuristic (line 9 in Algorithm 4.5) is labelled “Selection of neighbour” and all other routines are collectively labelled “Other”. The comparison in Figure 5.7 was generated using the Intel Xeon E7-8890 v3 running at 2.50 GHz with 144 cores. The Python profiler named “cProfile” was used to provide the performance breakdown.
@@
In Figure 5.7, we observe a clear trend in which the fraction of the computational time required for calculating gradient increases as the problem size increases. This increase in time is due primarily to the fact that these operations involve matrix–vector multiplications. As the problem size increases, the matrix–vector multiplications become more demanding due to the quadratic growth of the binary formulation. It is worth noting, however, that the relative cost of the gradient computation can vary depending on the computational resources used. Given that these steps are computational bottlenecks, IMC hardware presents a promising opportunity for considerable improvements in terms of computation time. Unlike conventional processors, which suffer from frequent memory access overhead, IMC architectures perform computations directly within memory. As shown in Figure 4.3, our approach is well-suited for this paradigm. Additionally, the energy efficiency of IMC hardware is a notable by-product, as it reduces the energy overhead of conventional digital computers.
@@ Appendix A Supplementary Proof
Appendix B Supplementary Numerical Results
We present the relative optimality gaps obtained using the approximate gradient method with the Top10, WalkQAP, and SA heuristics on selected QAPLIB instances. An Intel Xeon E7-8890 v3 CPU running at 2.50 GHz with 144 cores was used for the experiment. Figure B.1 presents the relative optimality gaps $\operatorname*{\text{rel\_gap}_{\text{min}}}(10^{5})$ for selected instances from the QAPLIB library. The results demonstrate that the Top10 and WalkQAP heuristics consistently outperform the SA heuristic, yielding significantly smaller gaps across most instances.
[…]